A car travels 6km east then turns south for another 5km. Find the magnitude and direction of the resultant displacement.


1
Expert's answer
2021-10-14T18:38:42-0400


The mafnitude of resultant displacement


"d=\sqrt{6^2+5^2}=7.8\;\rm km"

Direction

"\theta=\tan^{-1}\left(\frac{5}{6}\right)=40^{\circ}\; \rm S\; of \;E"